A continuous random variable has a density function f (x) = 2 (5-x)/5, where 2<x<3. Calculate the probability
1
Expert's answer
2020-11-30T09:42:37-0500

At first, we will check that "f(x)" it is a density. We calculate "\int_2^32\frac{5-x}{5}dx=(2x-\frac{x^2}5)|_2^3=6-\frac{9}{5}-4+\frac{4}{5}=1". Thus, it is indeed a probability density function. We denote by "X" a random variable that has a density "f(x)" . We calculate:

"P(X\leq a)=\left\{\begin{matrix}
 1,a\geq3, \\
 (2x-\frac{x^2}5)|_2^a=2a-\frac{a^2}{5}-\frac{16}5,\,\,\,2<a<3, \\
0,\quad a<2.
\end{matrix}\right."